George drove to town at 50 mph and rove back home at 40 mph. What is the distance to town if the whole trip took 9 hours?
Found 2 solutions by jorel1380, Alan3354:
Answer by jorel1380(3719)

x/50+x/40=9
4x+5x=1800
9x=1800
x=200
The distance to town is 200 miles..

Answer by Alan3354(69443)

Another approach:
The average speed for a round trip is similar to parallel resistors, but times 2.
Avg speed = 2*40*50/(40+50) = 4000/90 mph
r = 400/9 mph
t = 9 hours
d = (400/9)*9 = 400 miles round trip
--> 200 miles one way